OpenStudy (anonymous):

Help here please: Evaluate the following, give exact value when possible: a. sec(225°):

OpenStudy (anonymous):

use the sum/difference rules: 225 = 180 + 45 -15 = 30 - 45

OpenStudy (callisto):

a. sec(225°) = 1/cos225 = 1/cos (180 +45) = 1/ -cos45 = 1/(-sqrt2 /2) = 2/sqrt2 = sqrt2

OpenStudy (anonymous):

Ok. One sec guys, Im going to try to solve the second one by myself

OpenStudy (anonymous):

mine was just a suggestion...you could also do 45-60 = -15 if you're more comfortable with that...

OpenStudy (anonymous):

Ok, so first its sin(30-45)=sin 30 cos45 - cos 30 sin 45

OpenStudy (callisto):

yes :)

OpenStudy (anonymous):

simplify....

OpenStudy (anonymous):

Then, its (1/2)(1/sqrt(2))-(sqrt(3)/2)(1/sqrt(2))

OpenStudy (callisto):

Yes, what's next?!

OpenStudy (anonymous):

then, its 1-sqrt(3) / 2sqrt(2)?

OpenStudy (callisto):

yes :)

OpenStudy (anonymous):

Is it by any chance = sqrt(2) - sqrt(6)/4?

OpenStudy (callisto):

yes, when you do the rationalisation

OpenStudy (anonymous):

so its correct right?

OpenStudy (callisto):

I think so, @dpaInc what do you think?

OpenStudy (anonymous):

great job! :)

OpenStudy (anonymous):

Thanks @callisto, Thanks @dpainc, As usual you guys are awesome!

OpenStudy (anonymous):

Man, you guys teach so much better than my teachers :P!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

sin (75) would be sqrt(2) + sqrt(6)/4. right?

OpenStudy (callisto):

yes, according to my stupid calculator :P

OpenStudy (anonymous):

lol. Alright thank you:)

OpenStudy (callisto):

welcome :)